OpenStudy (anonymous):

factor and simplify: cos^2x-sin^2xcos^2x ?

OpenStudy (anonymous):

I think my first step is to take out a cos^2x so now it would be: cos^2x(-sin^2x+1)?

OpenStudy (agent0smith):

Yes, \[\Large \cos^2x(1-\sin^2x)\]You should recognize what's inside the brackets.

OpenStudy (anonymous):

so it'd be \[\cos ^{2} (\cos ^{2})\] ? @agent0smith which would give me a final of cos^4x?

OpenStudy (anonymous):

I think that's right.....

OpenStudy (agent0smith):

The medal was a hint :P